Use the following information to draw XYZ's yield curve, using:

- Average real risk free rate of return is constant over 30 years at 1%

- Inflation rate over 30 years is expected to be 4% in the first five years, 2% in year 6-15, and 1% in year 16-30

- DRP is constant at 1.5%

- LP is constant at 0.5%

- MRP derived from MRP=(t-1)*0.1%
Answer #1

According to the information given:

each data is calculated and YTM is equal to = real risk free rate + Inflation + DRP + LP + MRP

MRP for example is calculated as follows:

year 5 = (5 - 1 ) * 0.1% = 0.4%

Year 10 = (10 -1 ) * 0.1% = 0.9%

Year real inflation rate DRP LP MRP ytm
1 1% 4% 1.50% 0.50% 0 7.00%
2 1% 4% 1.50% 0.50% 0.001 7.10%
3 1% 4% 1.50% 0.50% 0.002 7.20%
4 1% 4% 1.50% 0.50% 0.003 7.30%
5 1% 4% 1.50% 0.50% 0.004 7.40%
6 1% 2% 1.50% 0.50% 0.005 5.50%
7 1% 2% 1.50% 0.50% 0.006 5.60%
8 1% 2% 1.50% 0.50% 0.007 5.70%
9 1% 2% 1.50% 0.50% 0.008 5.80%
10 1% 2% 1.50% 0.50% 0.009 5.90%
11 1% 2% 1.50% 0.50% 0.01 6.00%
12 1% 2% 1.50% 0.50% 0.011 6.10%
13 1% 2% 1.50% 0.50% 0.012 6.20%
14 1% 2% 1.50% 0.50% 0.013 6.30%
15 1% 2% 1.50% 0.50% 0.014 6.40%
16 1% 1% 1.50% 0.50% 0.015 5.50%
17 1% 1% 1.50% 0.50% 0.016 5.60%
18 1% 1% 1.50% 0.50% 0.017 5.70%
19 1% 1% 1.50% 0.50% 0.018 5.80%
20 1% 1% 1.50% 0.50% 0.019 5.90%
21 1% 1% 1.50% 0.50% 0.02 6.00%
22 1% 1% 1.50% 0.50% 0.021 6.10%
23 1% 1% 1.50% 0.50% 0.022 6.20%
24 1% 1% 1.50% 0.50% 0.023 6.30%
25 1% 1% 1.50% 0.50% 0.024 6.40%
26 1% 1% 1.50% 0.50% 0.025 6.50%
27 1% 1% 1.50% 0.50% 0.026 6.60%
28 1% 1% 1.50% 0.50% 0.027 6.70%
29 1% 1% 1.50% 0.50% 0.028 6.80%
30 1% 1% 1.50% 0.50% 0.029 6.90%

Only Column 1 and 7 is used to create the chart as follows:

Select column 1 and 2 and then go to insert chart, select scatter with smooth lines and marker.

ytm 8.00% 7.00% 5.00% 4.00% 3.00% 1.00% 30 35 20 10 1% 0

With axis marked.

SewAAAABJRU5ErkJggg==